OpenStudy (anonymous):

Find the equation for the parabola with focus at (3, -3) and vertex at (3, 1).

OpenStudy (anonymous):

Find the latus rectum = 4* a where a is distance between vertex and focus so a = 4, latus rectum = 16 since it is vertical downward parabola, eqn. will be (x-1)^2 = -16*(y-3)

OpenStudy (anonymous):

I need it in standard form

OpenStudy (anonymous):

if you need it in form y = ax^2 + bx + c just open and rearrange the eqn. i just wrote.... x^2 - 2x +1 = -16(y-3)